a 1582 kg car is moving at 72.9 kn/h down a highway when the driver sees a turtle in the road. not wanting to hit it, the driver slows the car down to 37.9 kn/h over a distance of 26.3 m.

what force (in newtons) is necessary to slow down the car?
Answer #1

Here first we will use

v​​​​​​2=u​​​​​​2+2as

v-final speed. ,. u-intial speed , a- acceleration , s - distance travelled

( 37.9×5/18)2=(72.9×5/18)2 +2a×26.3

a=5.69 m/s​​​​​s2 (deaccleration)

Force on car= mass× acceleration

Force=1582×5.69=8999.61 N